OpenStudy (anonymous):

Six is the geometric mean of 18 and what number? 4 3 2 9

hartnn (hartnn):

let that number be x so 6^2 =18 x can u find x now ?

OpenStudy (anonymous):

3 and 2 is the wrong answers

hartnn (hartnn):

6*6 =36 36/18 = 2 2 is the correct answer.

OpenStudy (anonymous):

well they said it wasnt

hartnn (hartnn):

then *they* were wrong. 2 is correct........

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):